Two investments earn an annual income of $478.40. One investment earns an annual simple interest rate of 6.5%, and the other investment earns an annual simple interest rate of 5.4%. The total amount invested is $7,800. How much is invested in each account?
Answer by mananth(16949)

Investment Part I 6.50% per annum ----x
Investment part II 5.40% per annum ----y
The sum of the investments is $7,800.00
The sum of individual interests = $478.40
x+y=7800 ------------------------1
6.50%x+ 5.40%y= $478.40
Multiply by 100
6.5x+5.4y= $47,840.00 --------2

Multiply (1) by -6.5
we get

-6.5x-6.5y= -50700.00

Add this to (2)

0x-1.1y =-$2,860.00

divide by-1.1

y = $2,600.00 investment at 5.40%
Balance $5,200.00 investment at 6.50%
CHECK
$5,200.00 @ 6.50% $338.00
$2,600.00 @ 5.40% $140.40
Total -------------------- $478.40